OpenStudy (anonymous):

3. The number of visitors to a small pumpkin patch that opened in 2003 increased each year as shown in the table. Predict the number of visitors in 2015, assuming the increase continues at this same rate. Year 2003 ,2004, 2005, 2006 ,2007, 2008 Visitors 310 332 355 380 406 434 please help! using the exponential growth formula y = b(1 + r)^t

OpenStudy (mathmath333):

y = b(1 + r)^t assume 2003 as zeroth year and 2004 as the first year then using formula 310 =b(1 + r)^0 310=b-------(1) 332 = b(1 + r)^1 332=b(1+r)--------(2) from that find value of r 2015 will be 11th year so after finding r calculate b(1+r)^(11) =310(1+r)^(11)

OpenStudy (mathmath333):

\[\Huge \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#E6E600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{o}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{s}}\color{#E6E600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{}}\end{array} \]

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@mathmath333 wouldn't it be with 12th year?

OpenStudy (mathmath333):

yes sry it will be 12th

OpenStudy (anonymous):

so....would the answer be 697?

OpenStudy (mathmath333):

what is r ?

OpenStudy (anonymous):

1.07%

OpenStudy (mathmath333):

wolfram is givind 705 as accurate http://www.wolframalpha.com/input/?i=%3D310%281%2B11%2F155%29%5E%2812%29